# ReDit: Reward Dithering for Improved LLM Policy Optimization

### v2

@@ Abstract
DeepSeek-R1 has successfully enhanced Large Language Model (LLM) reasoning capabilities through its rule-based reward system. While it’s a “perfect” reward system that effectively mitigates reward hacking, such reward functions are often discrete. Our experimental observations suggest that discrete rewards can lead to gradient anomaly, unstable optimization, and slow convergence. To address this issue, we propose ReDit (Reward Dithering), a method that dithers the discrete reward signal by adding simple random noise. With this perturbed reward, exploratory gradients are continuously provided throughout the learning process, enabling smoother gradient updates and accelerating convergence. The injected noise also introduces stochasticity into flat reward regions, encouraging the model to explore novel policies and escape local optima. Experiments across diverse tasks and different LLMs demonstrate the effectiveness and efficiency of ReDit. On average, ReDit achieves performance comparable to vanilla GRPO with only approximately 10% the training steps, and furthermore, still exhibits a 4% performance improvement over vanilla GRPO when trained for a similar duration. Visualizations confirm significant mitigation of gradient issues with ReDit. Moreover, theoretical analyses are provided to further validate these advantages.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in Large Language Model (LLM) development(AI@Meta, 2025; Anthropic, 2024; OpenAI et al., 2024). Initially, RL from human feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019) was employed to align pre-trained LLMs with human preferences (Lang et al., 2024; Ouyang et al., 2022). This typically involves training a separate reward model (RM) on human preference data (Kaufmann et al., 2024), which then guides the LLM policy optimization (Lambert, 2025). While effective, this approach introduces considerable training overhead (Cao et al., 2024b). Subsequently, methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) were developed, enabling LLMs to learn directly from preference data and thus bypassing explicit RM training. However, these methods still require extensive collection of high-quality preference data. For reasoning tasks such as mathematics and coding, DeepSeek-R1 (DeepSeek-AI et al., 2025) with Group Relative Policy Optimization (Shao et al., 2024)(GRPO) proposes an alternative: optimizing the LLM policy directly using a rule-based reward system (Kong & Yang, 2022; Wang et al., 2025), thereby avoiding the need for external RMs or large preference datasets. For instance, such a system might assign a reward of 1 for outputs meeting predefined criteria (e.g., correctness, format compliance) and 0 otherwise (DeepSeek-AI et al., 2025). The simplicity and unbiased nature of these rule-based rewards prevent LLMs from hacking them, potentially fostering enhanced reasoning capabilities (Chan et al., 2023).
Figure 1: Training Dynamics of Gradient Norm and Reward for Qwen2.5-7B (Qwen et al., 2025) on GSM8K Dataset. Fig.s (a) and (b) compare gradient distributions and reward trends across training steps. The original GRPO method (Fig. (a)) suffers from significant gradient instability—both vanishing (red dots, norms ¡ 0.01) and exploding (purple asterisks, norms ¿ 5). In contrast, ReDit with Gaussian reward smoothing (Fig. (b)) effectively stabilizes optimization throughout training.
@@
However, such reward functions are often discrete, posing significant optimization challenges (Rengarajan et al., 2022; Vasan et al., 2024; Goyal et al., 2019). Consider an RL scenario with a binary reward (Chatterji et al., 2021): a policy model receives 1 for a correct answer and 0 otherwise. During early training phases, a policy LLM rarely generates completely correct answers, resulting in predominantly zero rewards across mini-batches (Cao et al., 2024a). Although the model may engage in exploratory behavior on difficult examples, the corresponding gradients remain minimal due to small advantage magnitudes (Chan et al., 2024). Thus, these hard examples and potentially beneficial explorations (Chan et al., 2024) are largely unexploited during the early stages. Conversely, the model may repeatedly reinforce easy examples (Xie et al., 2024), thus reducing incentives to explore alternative strategies for more difficult problems (Weaver & Tao, 2001). This phenomenon can lead to training stagnation in intermediate and advanced stages. Consistent with this, as shown in Fig. 1(a), we observe that the policy model frequently suffers from gradient vanishing (Razin et al., 2024; Abdul Hameed et al., 2023) or explosion (Zhang et al., 2025) during these phases. This combination of insufficient exploration and gradient instability substantially impedes model convergence, representing a critical obstacle to efficient RL in LLM.
@@
Figure 2: The figure illustrates how ReDit of different variances gradually smooth the reward distribution, showing the smoothing effect of perturbations of different variances.
@@
This observed phenomenon highlights that even perfectly accurate discrete reward functions face significant limitations within gradient-based optimization frameworks. Lending theoretical support to this, recent studies (Ivison et al., 2024; Chen et al., 2024; Wen et al., 2025) have established that a singular focus on increasing reward model accuracy does not necessarily translate to enhanced language model performance. In particular, Wen et al. (2025) theoretically substantiates the necessity for effective reward models to integrate adequate variance and uncertainty to enable efficient optimization. The theoretical details are given in Sec. 3.2. Consequently, an optimal reward system must balance accuracy with appropriate levels of variance.
@@
Inspired by these observations and theoretical insights, we propose ReDit, a simple yet effective technique that applies zero-mean random perturbations to discrete reward signals during training. By introducing controlled noise to the reward function (Fig. 2), ReDit transforms hard reward boundaries into smoother gradients. This softened approach generates greater reward variance within mini-batches which, as established in previous research, enhances model performance and accelerates convergence.
@@
Fig.1 demonstrates the impact of ReDit on LLM policy optimization for GSM8K. The orange lines reveal that GRPO with ReDit achieves substantially higher early-phase rewards than the baseline, indicating the effectiveness of ReDit . We hypothesize that ReDit encourages broader exploration by assigning varied rewards to outputs that only partially meet strict evaluation criteria, thereby accelerating convergence. Although both approaches eventually attain high rewards after 1000 training steps, our method exhibits superior test set performance, suggesting better generalization. As evidenced by Fig.1(a), ReDit maintains stable gradients throughout training while the baseline suffers from gradient vanishing (red point) and explosion (purple star). These results confirm the advantages of ReDit: more stable optimization, faster convergence, and enhanced overall performance.
@@
Moreover, theoretical analysis indicates that a greater reward variance can enhance performance and accelerate convergence in RL (Wen et al., 2025). We increase reward variance within mini-batches while preserving the expected gradient through reward dithering. By carefully injecting noise into the reward function, ReDit achieves a balance between reward signal fidelity and reward variance, leading to enhanced policy optimization.
@@
  • Extensive experiments across diverse downstream tasks, RL algorithms, perturbation distributions and different LLMs demonstrate that ReDit achieves superior performance and enhanced convergence (Section 5).
  • Theoretical analysis proves that ReDit produces an unbiased estimate of the original gradient (Proposition 6.1), introduces beneficial gradient variance that mitigates vanishing and exploding gradients (Proposition 6.2), and significantly improves convergence speed (Proposition 6.3).
@@ 2 Preliminaries
We frame LLM generation as a sequential decision-making problem solvable via RL. The process is modeled as a Markov Decision Process (MDP) (Hallak et al., 2015) where the state $s_{t}=q;o_{<t}$ includes the prompt $q$ and generated tokens $o_{<t}$, the action $o_{t}$ is the next token selected from the vocabulary, and the policy $\pi_{\theta}(o_{t}|s_{t})$ is parameterized by $\theta$. The goal is to optimize the policy to maximize the expected sequence-level reward $R(q,o)=\sum_{t=1}^{|o|}r(s_{t},o_{t})$ over the prompt distribution $p_{Q}$:
@@
Recently, GRPO (Shao et al., 2024) was proposed as a PPO alternative that eliminates the need for independent RMs and value functions. GRPO typically processes sparse, discrete rewards directly, rather than continuous RM scores. For tasks like mathematical reasoning, this discrete reward $R(q,o)\in\{0,1\}$ is often determined by a simple function checking correctness or format. GRPO estimates the advantage $\hat{A}_{i,t}^{\text{GRPO}}$ by sampling $G$ responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their discrete rewards within the set. Its objective function, which includes a KL divergence term $D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ for stability, is given by:
$J_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim p_{Q}}\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\min\Bigl{(}r_{i,t}(\theta)\hat{A}_{i,t}^{\text{GRPO}},$ (2)
$\qquad\text{clip}\left(r_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}^{\text{GRPO}}\Bigr{)}\Biggr{]}$
$\quad-\beta\mathbb{E}_{q\sim p_{Q}}\left[D_{\text{KL}}(\pi_{\theta}(\cdot|q)||\pi_{\text{ref}}(\cdot|q))\right],$
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|s_{i,t})}$. Subsequent methods such as DAPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025), and REINFORCE++ (Hu et al., 2025) generally adopt this discrete reward paradigm (see Appendix A for more related work). While simplifying the overall RL process by avoiding complex RMs, this shift to discrete, sequence-level rewards introduces significant optimization challenges. The inherent sparsity and abrupt value changes (e.g., 0 to 1) hinder policy gradient estimation and lead to training instability (Section 3.1).
Figure 3: GRPO has unstable performance on the MATH test set. The figure plots the test accuracy achieved for the checkpoints saved during the training run shown in Fig. 4(b).
@@ 3.1 Difficulties in Optimization Caused by Discrete Rewards
Figure 4: Qwen2.5-7B Gradient norm and reward training dynamics of standard GRPO on GSM8k and MATH datasets. During the whole optimization process, the gradient of standard GRPO is unstable, and there are a lot of gradient vanishing or gradient exploding cases.
@@
Gradient Vanishing. The figure illustrates instances where the gradient norm approaches zero (red dot), occurring when most examples in a GRPO batch yield identical binary rewards. Consequently, the population relative advantage estimate $\hat{A}_{i,t}^{\text{GRPO}}$ becomes negligible across examples, providing insufficient learning signals and causing training stagnation. This phenomenon is evident in Fig. 4(b) post-step 2000.
Gradient Eplosion. Conversely, training dynamics exhibit sporadic sharp spikes in gradient norm (purple asterisks) when small policy changes cause sequences to transition from incorrect (reward 0) to correct (reward 1). These transitions create disproportionately large advantage estimates for newly successful sequences, triggering sudden, destabilizing gradient updates as shown in Fig. 4(a). Such spikes induce reward fluctuations in subsequent steps, hindering smooth convergence and learning efficiency.
@@ 3.2 Theoretical Principles to Address the Limitations of Discrete Rewards
To overcome the critical challenges with discrete rewards outlined in Section 3.1, we propose a approach to improve the quality of the reward signal. Our solution derives from Theorem 3.3 and Theorem 3.4, which reveal fundamental relationships between reward variance, accuracy, and learning efficiency. From Definition 1, 2 in Razin et al. (2025), The accuracy and variance of the reward function is as follows:
Definition 3.1.
$\text{acc}_{x,\mathcal{D}}(r_{RM}):=\mathbb{E}_{\{y,y^{\prime}\}\sim\mathcal{D}}\Biggl{[}\mathbbm{1}\biggl{[}\text{sign}\bigl{(}r_{RM}(x,y)$ (3)
$\qquad-r_{RM}(x,y^{\prime})\bigr{)}=\text{sign}\bigl{(}r_{G}(x,y)-r_{G}(x,y^{\prime})\bigr{)}\biggr{]}\Biggr{]},$
Definition 3.2.
$\text{Var}_{y\sim\pi_{\theta}(\cdot|x)}[r_{RM}$ $(x,y)]:=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\Biggl{[}\biggl{(}r_{RM}(x,y)$ (4)
$\qquad-\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot|x)}\bigl{[}r_{RM}(x,y^{\prime})\bigr{]}\biggr{)}^{2}\Biggr{]}.$
Theorem 3.3 establishes that the time $t_{\gamma}$required for policy improvement is inversely proportional to reward variance. When rewards exhibit insufficient variance—failing to adequately differentiate between high-quality and low-quality outputs under policy $\pi_{\theta}$, convergence slows significantly. This finding suggests that strategically increasing reward variance can accelerate policy convergence.
Theorem 3.3 (Policy network optimization time lower bound).
@@
From Theorem 1 in Razin et al. (2025). Suppose that we maximize the objective (Eq. (1)), using a general autoregressive policy $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})=\prod_{l=1}^{{\mathbf{y}}}\mathrm{softmax}({f_{\theta}({\mathbf{x}},{\mathbf{y}}_{<l})}_{{\mathbf{y}}_{l}})$. For any $\gamma>0$, prompt ${\mathbf{x}}\in{\mathcal{X}}$, and reward function $r$, the time it takes until $\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(t)}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}})]\geq\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}})]+\gamma$ is: $\Omega\left(\mathop{\mathbb{E}}\nolimits_{{\mathbf{x}}^{\prime}\sim S}\left[\mathrm{var}_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}}^{\prime})}(r({\mathbf{x}}^{\prime},{\mathbf{y}}))\right]^{-\frac{1}{3}}\right)$
Complementarily, Theorem 3.4 demonstrates that effective reward models must incorporate a calibrated degree of uncertainty. This controlled uncertainty creates essential exploration space during early training stages, preventing premature convergence and facilitating more efficient optimization.
Theorem 3.4 (Policy network optimization time upper bound).
@@
From Theorem 2 in Razin et al. (2025). Assume $\pi_{\theta}$ is a policy of the form $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})=\mathrm{softmax}[\theta_{:,{\mathbf{x}}}]_{{\mathbf{y}}}$. Given a hint ${\mathbf{x}}\in S$, let $\gamma>0$ and denote by $t_{\gamma}>0$ the initial time of $\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(t)}(\cdot|{\mathbf{x}})}[r_{\mathrm{G}}({\mathbf{x}},{\mathbf{y}})]\geq\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}})}[r_{\mathrm{G}}({\mathbf{x}},{\mathbf{y}})]+\gamma$. For any initial policy $\pi_{\theta(0)}$, a perfect RM converges to $t_{\gamma}$ that can be arbitrarily large, while a relatively inaccurate RM has an upper bound of ${\mathcal{O}}(\pi_{\theta(0)}(y^{\gamma}|{\mathbf{x}})^{-1})$.

### v4

@@ Abstract
DeepSeek-R1 has successfully enhanced Large Language Models (LLMs) reasoning capabilities through its rule-based reward system. While it’s a “perfect” reward system that effectively mitigates reward hacking, such reward functions are often discrete. Our experimental observations suggest that discrete rewards can lead to gradient anomaly, unstable optimization, and slow convergence. To address this issue, we propose ReDit (Reward Dithering), a method that dithers the discrete reward signal by adding simple random noise. With this perturbed reward, exploratory gradients are continuously provided throughout the learning process, enabling smoother gradient updates and accelerating convergence. The injected noise also introduces stochasticity into flat reward regions, encouraging the model to explore novel policies and escape local optima. Experiments across diverse tasks and different LLMs demonstrate the effectiveness and efficiency of ReDit. On average, ReDit achieves performance comparable to vanilla GRPO with only approximately 10% the training steps, and furthermore, still exhibits a 4% performance improvement over vanilla GRPO when trained for a similar duration. Visualizations confirm significant mitigation of gradient issues with ReDit. Moreover, theoretical analyses are provided to further validate these advantages.
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in Large Language Models (LLMs) development AIAtMeta2024Llama ; Anthropic2024Claude ; openai2024gpt4technicalreport ; wei-etal-2025-flexora . Initially, RL from human feedback (RLHF) DeepReinforcementLearningFromHumanPreferences ; ziegler2019finetuning was employed to align pre-trained LLMs with human preferences lang-etal-2024-fine ; ouyang2022training . This typically involves training a separate reward model (RM) on human preference data kaufmann2024surveyreinforcementlearninghuman , which then guides the LLM policy optimization lambert2025reinforcementlearninghumanfeedback . While effective, this approach introduces considerable training overhead Cao_2024 . Subsequently, methods like Direct Preference Optimization (DPO) rafailov2023direct ; wei2025testtimepolicyadaptationenhanced were developed, enabling LLMs to learn directly from preference data and thus bypassing explicit RM training. However, these methods still require extensive collection of high-quality preference data. For reasoning tasks such as mathematics and coding, DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability with Group Relative Policy Optimization shao2024deepseekmathpushinglimitsmathematical (GRPO) proposes an alternative: optimizing the LLM policy directly using a rule-based reward system ActiveRewardLearning ; wang2025crowdvlmr1expandingr1ability , thereby avoiding the need for external RMs or large preference datasets. For instance, such a system might assign a reward of 1 for outputs meeting predefined criteria (e.g., correctness, format compliance) and 0 otherwise deepseekai2025deepseekr1incentivizingreasoningcapability . The simplicity and unbiased nature of these rule-based rewards prevent LLMs from hacking them, potentially fostering enhanced reasoning capabilities chan2023visionlanguage .
@@
However, such reward functions are often discrete, posing significant optimization challenges rengarajan2022reinforcement ; vasan2024revisitingsparserewardsgoalreaching ; reward_shaping . Consider an RL scenario with a binary reward once-per-episodeFeedback : a policy model receives 1 for a correct answer and 0 otherwise. During early training phases, a policy LLM rarely generates completely correct answers, resulting in predominantly zero rewards across mini-batches cao-etal-2024-enhancing . Although the model may engage in exploratory behavior on difficult examples, the corresponding gradients remain minimal due to small advantage magnitudes chan2024dense . Thus, these hard examples and potentially beneficial explorations chan2024dense are largely unexploited during the early stages. Conversely, the model may repeatedly reinforce easy examples xie2024textreward , thus reducing incentives to explore alternative strategies for more difficult problems gradient-basedReinforcementLearning . This phenomenon can lead to training stagnation in intermediate and advanced stages. Consistent with this, as shown in Fig. 1 (the left figure), we observe that the policy model frequently suffers from gradient vanishing razin2024vanishing ; GradientMonitoredReinforcementLearning or explosion Evolution-guided ; zhang2025gvpogroupvariancepolicy during these phases. This combination of insufficient exploration and gradient instability substantially impedes model convergence, representing a critical obstacle to efficient RL in LLMs.
Figure 1: Training Dynamics of Gradient Norm and Reward for Qwen2.5-7B qwen2025qwen25technicalreport on GSM8K cobbe2021trainingverifierssolvemath Dataset. The left and right figures compare original gradient norm (before gradient clipping Zhang2020Why ) and reward trends across training steps. The original GRPO method (the left figure) suffers from significant gradient instability—both vanishing (red dots, norms < 0.01) and exploding (purple asterisks, norms > 5). In contrast, ReDit with Gaussian reward smoothing (the right figure) effectively stabilizes optimization throughout training.
@@
Figure 2: The figure illustrates how ReDit of different variances gradually smooth the reward distribution, showing the smoothing effect of perturbations of different variances on the reward distribution.
@@
This observed phenomenon highlights that even perfectly accurate discrete reward functions face significant limitations within gradient-based optimization frameworks. Lending theoretical support to this, recent studies ivison2024unpacking ; chen-etal-2024-accuracy ; wen2025rethinking have established that a singular focus on increasing reward model accuracy does not necessarily translate to enhanced language model performance. In particular, wen2025rethinking theoretically substantiates the necessity for effective reward models to integrate adequate variance and uncertainty to enable efficient optimization. The theoretical details are given in Sec. 3.2. Consequently, we believe that an excellent reward system should achieve a careful equilibrium between correctness and sufficient variance.
@@
Inspired by these observations and theoretical insights, we propose ReDit, a simple but effective technique that introduces zero-mean random perturbations to discrete reward signals during training. By adding small random noise to the reward function (Fig. 2), ReDit effectively softens the original hard reward function. Compared to a hard reward function, a softened one can provide greater reward variance within mini-batches, which, as indicated by previous research, can lead to enhanced model performance and accelerated convergence.
@@
Fig. 1 (the right figure) illustrates the impact of our proposed ReDit on LLM policy optimization for the GSM8K cobbe2021trainingverifierssolvemath task. The orange lines indicate that during early training phases, GRPO with ReDit achieves significantly higher average rewards compared to the baseline (GRPO without ReDit), demonstrating the efficacy of our method. We hypothesize that ReDit encourages broader exploration by assigning varied rewards to outputs that only partially meet strict evaluation criteria, thereby accelerating convergence. Towards the end of training (1000 steps), while both policy models attain high rewards, our approach with ReDit exhibits superior performance on the test set, indicating enhanced generalization. Additionally, ReDit demonstrates more robust gradient updating. As shown in Fig. 1 (the left figure), phenomena such as gradient vanishing (red point) and explosion (purple star) emerge during training with the baseline. In contrast, GRPO with ReDit maintains stable gradients throughout the training process. These findings highlight the advantages of ReDit: more stable policy optimization, faster convergence, and improved overall performance.
@@
Moreover, theoretical analysis indicates that a greater reward variance can enhance performance and accelerate convergence in reinforcement learning wen2025rethinking . We increase reward variance within mini-batches while preserving the expected gradient through reward dithering. By carefully injecting noise into the reward function, ReDit achieves a balance between reward signal fidelity and reward variance, leading to enhanced policy optimization.
@@
  • Extensive experiments across diverse downstream tasks, reinforcement learning algorithms, and perturbation distributions demonstrate that ReDit achieves superior performance and enhanced convergence (Section 5).
  • Theoretical analysis proves that ReDit produces an unbiased estimate of the original gradient (Proposition 1) and introduces beneficial gradient variance that mitigates vanishing and exploding gradients (Proposition 2).
@@ 2 Preliminaries
We frame LLM generation as a sequential decision-making problem solvable via RL. The process is modeled as a Markov Decision Process (MDP) hallak2015contextualmarkovdecisionprocesses where the state $s_{t}=q;o_{<t}$ includes the prompt $q$ and generated tokens $o_{<t}$, the action $o_{t}$ is the next token selected from the vocabulary, and the policy $\pi_{\theta}(o_{t}|s_{t})$ is parameterized by $\theta$. The goal is to optimize the policy to maximize the expected sequence-level reward $R(q,o)=\sum_{t=1}^{|o|}r(s_{t},o_{t})$ over the prompt distribution $p_{Q}$:
@@
Recently, GRPO shao2024deepseekmathpushinglimitsmathematical was proposed as a PPO alternative that eliminates the need for independent RMs and value functions. GRPO typically processes sparse, discrete rewards directly, rather than continuous RM scores. For tasks like mathematical reasoning, this discrete reward $R(q,o)\in\{0,1\}$ is often determined by a simple function checking correctness or format. GRPO estimates the advantage $\hat{A}_{i,t}^{\text{GRPO}}$ by sampling $G$ responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their discrete rewards within the set. Its objective function, which includes a KL divergence term $D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ for stability, is given by:
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim p_{Q}}\left[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\min\left(r_{i,t}(\theta)\hat{A}_{i,t}^{\text{GRPO}},\text{clip}\left(r_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}^{\text{GRPO}}\right)\right]$ (2)
$-\beta\mathbb{E}_{q\sim p_{Q}}\left[D_{\text{KL}}(\pi_{\theta}(\cdot|q)||\pi_{\text{ref}}(\cdot|q))\right],$
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|s_{i,t})}$. Subsequent methods such as DAPO yu2025dapoopensourcellmreinforcement , Dr.GRPO liu2025understandingr1zeroliketrainingcritical , and REINFORCE++ hu2025reinforceefficientrlhfalgorithm generally adopt this discrete reward paradigm (see Appendix A for more related work). While simplifying the overall RL process by avoiding complex RMs, this shift to discrete, sequence-level rewards introduces significant optimization challenges. The inherent sparsity and abrupt value changes (e.g., 0 to 1) hinder policy gradient estimation and lead to training instability (Section 3.1).
@@ 3.1 Difficulties in Optimization Caused by Discrete Rewards
Figure 3: Qwen2.5-7B qwen2025qwen25technicalreport Gradient norm and reward training dynamics of standard GRPO on GSM8k and MATH datasets. During the whole optimization process, the gradient of standard GRPO is unstable, and there are a lot of gradient vanishing or gradient exploding cases.
@@
Figure 4: GRPO has unstable performance on the MATH test set. The figure plots the test accuracy achieved for the checkpoints saved during the training run shown in Fig. 3(the right figure).
Gradient Vanishing. The figure illustrates instances where the gradient norm approaches zero (red dot), occurring when most examples in a GRPO batch yield identical binary rewards. Consequently, the population relative advantage estimate $\hat{A}_{i,t}^{\text{GRPO}}$ becomes negligible across examples, providing insufficient learning signals and causing training stagnation. This phenomenon is evident in Fig. 3(the right figure) post-step 2000.
Gradient Explosion. Conversely, training dynamics exhibit sporadic sharp spikes in gradient norm (purple asterisks) when small policy changes cause sequences to transition from incorrect (reward 0) to correct (reward 1). These transitions create disproportionately large advantage estimates for newly successful sequences, triggering sudden, destabilizing gradient updates as shown in Fig. 3(the left figure). Such spikes induce reward fluctuations in subsequent steps, hindering smooth convergence and learning efficiency.
@@ 3.2 Theoretical Principles to Address the Limitations of Discrete Rewards
To overcome the critical challenges with discrete rewards outlined in Section 3.1, we propose a direct approach to enhance reward signal quality. Our solution derives from key theoretical insights in policy optimization, particularly Theorem 1 and Theorem 2, which reveal fundamental relationships between reward variance, model accuracy, and learning efficiency. See Appendix B.1 for definitions.
Theorem 1 (Policy network optimization time lower bound).
@@
From Theorem 1 in razin2025makesrewardmodelgood . Suppose that we maximize the objective (Eq. (1)), using a general autoregressive policy $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})=\prod_{l=1}^{{\mathbf{y}}}\mathrm{softmax}({f_{\theta}({\mathbf{x}},{\mathbf{y}}_{<l})}_{{\mathbf{y}}_{l}})$. For any $\gamma>0$, prompt ${\mathbf{x}}\in{\mathcal{X}}$, and reward function $r$, the time it takes until $\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(t)}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}})]\geq\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}})]+\gamma$ is: $\Omega\left(\mathop{\mathbb{E}}\nolimits_{{\mathbf{x}}^{\prime}\sim S}\left[\mathrm{var}_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}}^{\prime})}(r({\mathbf{x}}^{\prime},{\mathbf{y}}))\right]^{-\frac{1}{3}}\right)$ (3) The reward variance is: $\mathrm{var}_{{\mathbf{y}}\sim\pi_{\theta}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}})]:=\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta}(\cdot|{\mathbf{x}})}[[r({\mathbf{x}},{\mathbf{y}})-\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}^{\prime}\sim\pi_{\theta}(\cdot|{\mathbf{x}})}[r({\mathbf{x}},{\mathbf{y}}^{\prime})]]^{2}]$.
Theorem 1 establishes that the time $t_{\gamma}$required for policy improvement is inversely proportional to reward variance. When rewards exhibit insufficient variance—failing to adequately differentiate between high-quality and low-quality outputs under policy $\pi_{\theta}$, convergence slows significantly. This finding suggests that strategically increasing reward variance can accelerate policy convergence.
Theorem 2 (Policy network optimization time upper bound).
@@
From Theorem 2 in razin2025makesrewardmodelgood . Assume $\pi_{\theta}$ is a policy of the form $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})=\mathrm{softmax}[\theta_{:,{\mathbf{x}}}]_{{\mathbf{y}}}$. Given a prompt ${\mathbf{x}}\in S$, let $\gamma>0$ and denote by $t_{\gamma}>0$ the initial time of $\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(t)}(\cdot|{\mathbf{x}})}[r_{\mathrm{G}}({\mathbf{x}},{\mathbf{y}})]\geq\mathop{\mathbb{E}}\nolimits_{{\mathbf{y}}\sim\pi_{\theta(0)}(\cdot|{\mathbf{x}})}[r_{\mathrm{G}}({\mathbf{x}},{\mathbf{y}})]+\gamma$. For any initial policy $\pi_{\theta(0)}$, a perfect RM converges to $t_{\gamma}$ that can be arbitrarily large, while a relatively inaccurate RM has an upper bound of ${\mathcal{O}}(\pi_{\theta(0)}(y^{\gamma}|{\mathbf{x}})^{-1})$.
Complementarily, Theorem 2 demonstrates that effective reward models must incorporate a calibrated degree of uncertainty. This controlled uncertainty creates essential exploration space during early training stages, preventing premature convergence and facilitating more efficient optimization.
@@
4 Reward Dithering (ReDit)
The crucial modification introduced by ReDit lies in how the advantage term $\hat{A}_{i,t}^{\text{GRPO}}$ within Eq. (2) is computed. Instead of directly using the raw discrete rewards $r_{i}=r(o_{i})$ obtained for each sampled output $o_{i}$ in the batch $\{o_{i}\}_{i=1}^{G}$ (line 3 in Algorithm 1), we first compute smoothed rewards $\tilde{r}_{i}$. This is done by adding independently sampled zero-mean perturbation $\epsilon_{i}$ (e.g., from $\mathcal{N}(0,\sigma^{2})$ or $\mathcal{U}[-a,a]$) to each discrete reward (line 4 in Algorithm 1):
$\tilde{r}_{i}=r_{i}+\epsilon_{i}$ (4)
$\hat{A}_{i,t}^{\text{GRPO}}\propto\frac{r_{i}-\text{mean}(\{{r}_{k}\}_{k=1}^{G})}{\text{std}(\{{r}_{k}\}_{k=1}^{G})}\quad\xrightarrow{\text{{ReDit}{}}}\quad\hat{A}_{i,t}^{\text{Dithering}}\propto\frac{\tilde{r}_{i}-\text{mean}(\{\tilde{r}_{k}\}_{k=1}^{G})}{\text{std}(\{\tilde{r}_{k}\}_{k=1}^{G})}$ (5)
Thus, the relative standing of each output $o_{i}$ within the batch, which informs its advantage $\hat{A}_{i,t}^{\text{GRPO}}$ used in Eq. (2), is determined by the continuous smoothed reward $\tilde{r}_{i}$ rather than the discrete $r_{i}$. This substitution transforms the optimization landscape. By introducing continuous variations via $\tilde{r}_{i}$, the added noise provides informative, non-zero gradients even when discrete rewards $r_{i}$ are sparse or identical within a batch, mitigating gradient vanishing. It also dampens the sharp changes in expected advantage resulting from small policy shifts affecting discrete outcomes, thus reducing the likelihood of gradient explosion. This overall smoothing effect facilitates a more stable gradient flow, enabling more robust and efficient optimization of the policy $\pi_{\theta}$ using the GRPO objective (line 5 in Algorithm 1).